OpenStudy (anonymous):

Two non-simultaneous events have an equal likelihood of occurring (50%). If the events are independent of each other, what is the probability that they will both occur? 30% 25% 50% 35% 40%

OpenStudy (anonymous):

help me @aaronq

OpenStudy (anonymous):

@geerky42

geerky42 (geerky42):

Sorta like asking the probability of getting head twice after flipping coin twice.

geerky42 (geerky42):

So when you flip coin twice, what is probability that you will get head twice?

geerky42 (geerky42):

Pretty much just \(\dfrac{1}{2}\cdot\dfrac{1}{2}\)

OpenStudy (anonymous):

what 100

geerky42 (geerky42):

Huh?

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

so it is 50

geerky42 (geerky42):

1/2*1/2 = 50%?

geerky42 (geerky42):

I thought \(\dfrac{1}{2}\times\dfrac{1}{2} = \dfrac{1}{4}\)?

geerky42 (geerky42):

What is 1/4 in percentage?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

@danieke

geerky42 (geerky42):

Yep. 25%

OpenStudy (anonymous):

thanks you